OpenStudy (anonymous):

HELP PLZZZ f(x) = 1/x determinate what is f(f(x)) and its domain

OpenStudy (anonymous):

try this domain of f is the values of f(x) for which the fuction is defined. in this case. (3 - x) > = 0 ( square root has to be positive) so, 3 > =x so, x<=3. so, domain = ( - infinity, 3]

OpenStudy (lgbasallote):

f(x)ception

OpenStudy (anonymous):

f(f(x)) = f(1/x) = 1/(1/x) = x Domain = R

OpenStudy (anonymous):

isn't the domain R except 0 ?

OpenStudy (anonymous):

f(f(x)) = x What do you think?

OpenStudy (anonymous):

welll, f(f(x)) = 1/(1/x). This is the original form, thus x can't be 0

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

1/(1/x) = x

OpenStudy (anonymous):

x not equal 0

OpenStudy (anonymous):

Domain = R-{0}

OpenStudy (anonymous):

p/q q not equal 0 remember

OpenStudy (anonymous):

lol can anyone confirm whatever 0 is included or excluded? :S its confusing

OpenStudy (amistre64):

the limit as x -> 0 of f(f(x)) = 1 ;) but yes, as is, the domain is restricted by the parts that make it up, not by the algebra that defines an equivalent curve

OpenStudy (anonymous):

so 0 is excluded??

ganeshie8 (ganeshie8):

see if this helps : domain of f(f(x)) = range of f(x) whats the range of f(x) ?

OpenStudy (amistre64):

1/x excludes the zero

OpenStudy (anonymous):

Well explained @ganeshie8.

OpenStudy (anonymous):

OK

OpenStudy (amistre64):

why did i type "1" ?? the limit as x -> 0 of f(f(x)) = 0

OpenStudy (anonymous):

the domain is R

OpenStudy (amistre64):

abay, no it isnt

OpenStudy (amistre64):

ganesh, are you reading that as tho its an inverse? or am i misreading it?

OpenStudy (amistre64):

i see it now ... whew!! thought i was going senile lol

OpenStudy (anonymous):

thx guys!